a)You want to prepare 1.00 L of a buffer solution with a pH of 3.5. You are given an 0.45 M solution of nitrous acid, HNO2, and an 0.68 M solution of NaOH. What volume of each of these solutions must be mixed in order to get the desired buffer solution? Use the simultaneous equations method to solve this problem.

b)If the acid used in a. was nitric acid, HNO3, could the same buffer solution be made? Briefly explain.

c)If the buffer solution prepared in a. also contained ~10^–6M Ni^2+ and ~10^–6 M Pb^2+ , could those metal ions be separated from each other by saturating the solution with H2S? Explain your answer by showing the necessary calculations.
1
Expert's answer
2020-12-11T02:50:01-0500

a) If we give NaOH 100 ml. pKa=3.39


pH=pKa+log([A−]/[HA]).


log([A−]/[HA]) = pH-pKa.


log([A−]/[HA])= 3.5-3.39 = 0.11.


[A−]= 0.1 x 0.68 = 0.068.


log(0.68/[HA])=0.11.


[HA]= 0.071.


0.071 + 0.068 = 0.139.


0.139/0.45 = 0.3088 L. = 308.8 ml.


b)


b) No, because Buffer Solution is a water solvent based solution which consists of a mixture containing a weak acid and the conjugate base of the weak acid, or a weak base and the conjugate acid of the weak base. 


c) No, because there NaNO2.